OpenStudy (vshiroky):

Write an equation of the line passing through the given points. (4,-18) and (-1,7) Can someone please do this one for me? I tried I'm lost and have to leave for work very soon.

OpenStudy (jdoe0001):

hmmm haven't you done about 20 like this already?

OpenStudy (vshiroky):

Yes but I get confused when it ends up with a fraction

OpenStudy (vshiroky):

I got y+18=25/3(x-4)

OpenStudy (vshiroky):

That is as far as I can get

OpenStudy (johnweldon1993):

so slope would be 7 + 18 25 ------ = ----- = -5 -1 - 4 -5 Now you use that and 1 of your points to solve your equation for 'b' y = mx + b lets use (-1,7) and m = -5 7 = -5(-1) + b 7 = 5 + b b = 2 So your equation is y = -5x + 2

OpenStudy (vshiroky):

oh gosh now that I see it typed up I see I made a typo in my equation when I was trying to do it

OpenStudy (johnweldon1993):

mmhmm just a little mistake can do it :)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ -18}})\quad &({\color{red}{ -1}}\quad ,&{\color{blue}{ 7}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ 7}}-{\color{blue}{ -18}}}{{\color{red}{ -1}}-{\color{red}{ 4}}}\implies \cfrac{7+18}{-5}\implies \cfrac{25}{-5}\implies -5 \\ \quad \\ \)